Antonio has 5 times as much money as Eric. If Antonio gives 640 to Eric, they will have the same amount of money. HOw much money does Antonio have at first?
Answer by Cromlix(4381)

Antonio has 5 times as much money as Eric.
Eric has x
Antonio has 5x
If Antonio gives 640 to Eric, they will have the same amount of money.
5x - 640 = x + 640
Collect like terms:
5x - x = 640 + 640
4x = 1280
x = 320.
At start:
Eric had 320
Antonio had 1600.
Hope this helps.
:-)